OpenStudy (anonymous):

solve {2x-3y=11 7x-4y=6

OpenStudy (anonymous):

7*(2x-3y=11)= 11 2*(7x-4y=6)=6 multiplying both sides 13y=65 so y=5 so calculate x

OpenStudy (anonymous):

sry.. copy and paste made sm typo... but answer will be same 7*(2x-3y=11) 2*(7x-4y=6) multiplying both sides 13y=65 so y=5 so calculate x

OpenStudy (anonymous):

it will cm to 14x - 21y= 77 14x-18y=12 - + - (reversing the signs) -------------------------------- 13y=65 y=5

OpenStudy (anonymous):

is that matches ur answer Ashley?

OpenStudy (anonymous):

I appreciate ur help but i need to come out with an ordered pair for a graph. my answer sheet says the answer is (-2, -5)

OpenStudy (anonymous):

umm I think amit7808 made an error... I get your result lemme write the way I calculated it down:

OpenStudy (anonymous):

2x-3y=11 multiply by 4 since it is 4y in the equation below and we want to eliminate y 7x-4y=6 multiply by 3 since it is 3y in equation above so we can subtract both to obtain a solution for x => 8x-12y=44 21x+12y=18 :now we subtract the bottom one from the top and have a new bottom equation => 8x-12y=44 :top remains the same -13x =26 :bottom after subtracting, now we divide both sides by -13 so that we are left with 1x since -13/-13=1 and 1x=x on the left giving us the solution to x => x=26/-13= -2 we now insert -2 for x into the top equation to obtain y => 8*(-2)-12y=44 (one could also take the original version: 2*(-2)-3y=11) <=> -16-12y=44 now add 16 to both sides -12y=60 now same as before divide by minus twelve on both sides so we get 1*y=y => y=60/-12=-5 hope it is clear enough

OpenStudy (anonymous):

ya it will be -13y=65 so, y=-5 and x=-2 sry..